A stream function is given by Ψ = 2x- 5y. Calculate the velocity components




and also magnitude and direction of the resultant velocity at any point

1
Expert's answer
2022-08-10T16:04:07-0400

In velocity we can write


"u=\dfrac{\partial\psi}{\partial y}, v=-\dfrac{\partial\psi}{\partial x}"

Given "\psi=2x-5y"

"u=\dfrac{\partial\psi}{\partial y}=-5, v=-\dfrac{\partial\psi}{\partial x}=-2""u=-5, v=-2""|velocity|=\sqrt{u^2+v^2 }=\sqrt{(-5)^2+(-2)^2}=\sqrt{29}""\tan\theta=\dfrac{v}{u}=\dfrac{-2}{-5}=0.4"

The vector of the resultant velocity at any point consists the angle "\tan^{-1}(0.4)\approx21.8\degree" with the positive direction of the x-axis.